help me please

After flying for 14 min in a wind blowing 36 km/h
Answer #1

refer to the attached figure.

t = 14 m = 0.233 hrs

the velocity of plane relative to ground will be:

v = 56/0.233 = ( 240.34 ) j

and v' = 36 ( cos16 i - sin16 j) = 34.61 i - 9.22 j

v = V + v' => V = v - v'

V = 240.34 j - (34.61 i - 9.22 j) = -34.61 i + 249.56 j

V = sqrt [ (34.61)^2 + (249.56)^2 ] = 251.95 km/h

Hence, V = 251.95 km/h = 252 km/h.